OpenStudy (anonymous):

I can't understand why, after evaluating this, I get √3/2. I need to evaluate cos(sin^-1(-3/5). sin^-1 is the Inverse of sin, or arcsin. I end up with cos(4/5), after performing the pythagorean theorem getting x=4. When I draw out cos(4/5) "x/r" with -3 as y, I end up with √3/2 with my knowledge of right triangles on the unit circle. I think I am wrong here, would anyone else evaluate this a different way?

OpenStudy (anonymous):

cos(sin^-1(-3/5) Where is the missing parentheses ? There are two open and on closed.

OpenStudy (anonymous):

Sorry, "There are two open and one closed."

OpenStudy (anonymous):

\[\cos \left(\sin ^{-1}\left(-\frac{3}{5}\right)\right)=\frac{4}{5} \]

OpenStudy (anonymous):

\[\sqrt{25-9}=4 \]

OpenStudy (anonymous):

In a 5,4,3 right triangle one of the Cosines is 4/5.

OpenStudy (anonymous):

My calculator was terribly faulty. I was being lazy. Thank you. It just had me really confused.